CONSTRUCTION OF A RAMAN TWEEZER SYSTEM FOR INVESTIGATION OF BIOLOGICAL MOLECULES^{a}

\underline{GUNAY BASAR}, Technical University of Istanbul, Faculty of Science and Letters, Physics Department, 80626 Maslak, Istanbul, Turkey ; SEDA KIN, Technical University of Istanbul, Faculty of Science and Letters, Physics Department, 80626 Maslak, Istanbul, Turkey ; and SEVIM AKYUZ, Istanbul University, Faculty of Science, Physics Department, 34459 Vezneciler, Istanbul, Turkey.

Raman spectroscopy has been widely used in providing detailed analysis of a variety of biological samples. The combination of laser tweezer^{b} with Raman spectroscopy has allowed to get the Raman spectrum of a single living cell without causing any damage within their natural environment ^{c}. We have constructed a Raman tweezer system. In our set-up a temperature and current controlled 785 nm diode laser was employed for tweezing of micro particles in an aqua solution by using a high NA microscope objective. The same laser was also used for Raman excitation. We present here preliminary results of Raman spectra of some biological molecules and polystyrene micro spheres as test objects taken with our experimental set-up.

aSupport by the Scientific and Technological Research Council of Turkey under the project number 105T482 is gratefully acknowledged.

bA. Ashkin, J. M. Dziedzic, J. E. Bjorkholm and S. Chu Opt. Lett.\underline{11}, 288-300 1986.

cC. Xie, M. A. Dinno and Y. Li Opt. Lett.\underline{27}, 249-251 2002.